Procter & Gamble Is a multnational corporation that manufactures and markets many products that you use every day. In 2013, sales for the company were $91.500 (all amounts In milions). The annual report did not report the amount of credit sales, so we will assume that all sales were on credit. The average gross proflt percentage was 50.9 percent Account balances for that year tollow: Beginning Ending $ 5.300 6,000 6380 ,400 Accounts receivable (net) Required: 1. Compute the following tumover ratios. (Do not round intermediate calculations. Round your final answers to 1 decimal place.) Turnover Ratio Inventory Turnover R 2. By dividing 365 by your ratios from requirement 1, calculate the average days to collect receivables and the average days to sell inventory. (Round your intermediate calculations and final answers to 1 decimal place.) Average Days to Collect Receivables Aversge Days to Sell Inventory days days
Answer #1

The receivable turnover Ratio is an accounting measure used to quantify a firm's effectiveness in extending credit and in collecting debts on that credit.

Receivables turnover Ratio can be calculate by dividing the net credit sales during the given period by the average accounts receivable during the same period. Average accounts receivable can be calculate by adding the value of accounts receivable at the beginning of the desired period to their value at the end of the period and dividing the sum by two

Accounts Receivable Turnover= Net Credit sales/ Average Accounts Receivable

All amounts in millions

Therefore from the above Net credit sales are $91,500

Average Accounts Receivable= ($5600+$6000)/2=$5800

Turnover Ratio= $91,500/$5,800=15.77 times on average per year

Calculation of average days to collect Receivable= 365/15.77 =23.14 days

The average customer takes 23 days to pay his or her bills

Inventory turnover details how much inventory is sold over a period of time. To calculate inventory turnover Ratio cost of goods sold or sales is dividing the average inventory for the same period

Average inventory= (opening inventory+ closing inventory)/2

Average inventory= ($6380+$6400)/2=$6390

Inventory Turnover Ratio= $91,500/$6390=14.31

Average days to sell inventory= 365/14.31=25.50

This indicates company sells it's entire inventory within 26 days period